# Fractional Sobolev logarithmic inequalities

### v2

@@ 1. Introduction
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{p^{2}}\log\left[C_{p}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx\right],$ (1.1)
@@
$C_{p}=\frac{p}{d}\left(\frac{p-1}{e}\right)^{p-1}\pi^{-\frac{p}{2}}\left[\frac{\Gamma\left(\frac{d}{2}+1\right)}{\Gamma\left(d\frac{p-1}{p}+1\right)}\right]^{\frac{p}{d}}$ (1.2)
@@
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{sp^{2}}\log\left(\mathcal{C}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy\right),$ (1.3)
@@
$\mathcal{C}=C(d,p)\left(\frac{s(1-s)}{(d-sp)^{p-1}}\right),$ (1.4)
@@
$\lim_{s\to 1}(1-s)\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy=K\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx.$ (1.5)
@@
A version of the above theorem on stratified Lie groups was proved by Ghosh, Kumar, and Ruzhansky in [15, Section 7]. They established subelliptic fractional logarithmic Sobolev inequalities on stratified Lie groups with explicit constants depending on an additional parameter lying between $p$ and the fractional critical Sobolev exponent, under the condition $sp<d$. In contrast, Theorem 1.1 provides a constant $\mathcal{C}$ in Euclidean space that is independent of any extra parameter and also recovers the classical Sobolev logarithmic inequality.
@@
$\int_{\mathbb{R}^{d}}|u(x)|^{2}\log\left(\frac{|u(x)|^{2}}{\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}}\right)\,dx+d(1+\log c)\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}\leq\frac{c^{2}}{\pi}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{2}\,dx.$ (1.6)
@@
$\leq\frac{de^{p-1}c^{p}\,\mathcal{C}}{sp}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy,$ (1.7)
@@
$\leq\frac{de^{p-1}c^{p}\,K}{p}\left(\frac{C(d,p)}{(d-p)^{p-1}}\right)\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx.$ (1.8)
@@
$\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{(s-\alpha)p^{2}}\log\left(\mathcal{C}_{\alpha}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy\right),$ (1.9)
@@
$\mathcal{D}^{p,q}_{s}=(W^{s,p}(\mathbb{R}^{d})\cap L^{q}(\mathbb{R}^{d}))\setminus\{0\}.$ (1.10)
@@
$\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathscr{L}\,[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\,\|u\|^{1-a}_{L^{q}(\mathbb{R}^{d})},$ (1.11)
@@
$a=\frac{d(q-p)}{(q-1)\bigl(dp-(d-sp)q\bigr)},$ (1.12)
@@ 2. Proof of fractional Sobolev logarithmic inequalities
$\left(\int_{\mathbb{R}^{d}}|u(x)|^{p^{*}_{s}}\,dx\right)^{\frac{1}{p^{*}_{s}}}\leq\mathcal{C}^{\frac{1}{p}}\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy\right)^{\frac{1}{p}},$ (2.1)
@@
$\mathcal{C}=C(d,p)\left(\frac{s(1-s)}{(d-sp)^{p-1}}\right),$ (2.2)
@@
$\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathcal{C}^{\frac{a}{p}}[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a},$ (2.3)
@@
$\int_{\mathbb{R}^{d}}|u(x)|^{q}\log\left(\frac{|u(x)|}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)\,dx\leq\frac{r}{r-q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$ (2.5)
@@
$\log\|u\|_{L^{s}(\mathbb{R}^{d})}\leq\alpha(s)\log\|u\|_{L^{q}(\mathbb{R}^{d})}+(1-\alpha(s))\log\|u\|_{L^{r}(\mathbb{R}^{d})}.$ (2.6)
@@
$s\Phi^{\prime}(s)=\frac{1}{\|u\|^{s}_{L^{s}(\mathbb{R}^{d})}}\int_{\mathbb{R}^{d}}|u(x)|^{s}\log\left(\frac{|u(x)|}{\|u\|_{L^{s}(\mathbb{R}^{d})}}\right)\,dx.$ (2.7)
@@
$=\left(\frac{qr}{s^{2}(r-q)}\right)\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$ (2.8)
@@
$\frac{1}{a}\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)-\log\left(\mathcal{C}^{\frac{1}{p}}\right)\leq\log\left(\frac{[u]_{W^{s,p}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$ (2.9)
@@
$\frac{sp}{d}\int_{\mathbb{R}^{d}}\frac{|u(x)|^{p}}{\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}}\log\left(\frac{|u(x)|}{\|u\|_{L^{p}(\mathbb{R}^{d})}}\right)\,dx-\log\left(\mathcal{C}^{\frac{1}{p}}\right)\leq\log\left(\frac{[u]_{W^{s,p}(\mathbb{R}^{d})}}{\|u\|_{L^{p}(\mathbb{R}^{d})}}\right).$ (2.10)
@@ 3. Proof of Weighted fractional Sobolev logarithmic inequality
$\|u\|_{W^{s,p,\alpha}(\mathbb{R}^{d})}=\left(\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}+[u]^{p}_{W^{s,p,\alpha}(\mathbb{R}^{d})}\right)^{\frac{1}{p}}.$ (3.1)
@@
$\||x|^{\gamma}u\|_{L^{\tau}(\mathbb{R}^{d})}\leq C[u]^{a}_{W^{s,p,\alpha}(\mathbb{R}^{d})}\||x|^{\beta}u\|^{(1-a)}_{L^{q}(\mathbb{R}^{d})},\hskip 8.5359pt\forall\ u\in C^{1}_{c}(\mathbb{R}^{d}).$ (3.2)
@@
$\left(\int_{\mathbb{R}^{d}}|u(x)|^{\tau}\,dx\right)^{\!\frac{1}{\tau}}\leq C\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy\right)^{\!\frac{1}{p}},$ (3.3)
@@
$\forall\ u\in W^{s,p,\alpha}(\mathbb{R}^{d}).$ (3.4)
@@
$\mathcal{D}^{p,q}_{s,\alpha}=(W^{s,p,\alpha}(\mathbb{R}^{d})\cap L^{q}(\mathbb{R}^{d}))\setminus\{0\}.$ (3.5)
@@
$\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathcal{C}_{\alpha}^{\frac{a}{p}}[u]_{W^{s,p,\alpha}(\mathbb{R}^{d})}^{a}\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a},$ (3.6)
@@
$\leq\frac{d\,e^{p-1}c^{p}\,\mathcal{C}_{\alpha}}{(s-\alpha)p}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy.$ (3.8)
@@
$\inf_{s>0}f(s)=f\left(\left(\frac{bN}{aM}\right)^{\frac{1}{a+b}}\right)=\frac{a+b}{a}\left(\frac{b}{a}\right)^{\frac{-b}{a+b}}M^{\frac{b}{a+b}}N^{\frac{a}{a+b}}.$ (4.1)
@@
$I(u):=\frac{1}{p}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy+\frac{1}{q}\int_{\mathbb{R}^{d}}|u(x)|^{q}\,dx.$ (4.2)
@@
Now set $a=\alpha p-(d-sp)>0$, $b=d-\alpha q>0$, $M=\frac{1}{p}[u]^{p}_{W^{s,p}(\mathbb{R}^{d})}>0$, and $N=\frac{1}{q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}>0$. Applying Lemma 4.1 with these parameters, we obtain
$\inf_{u\in\Lambda}I(u)$ $=\inf_{u\in\Lambda}I(u)\,\inf_{\lambda>0}I(u_{\lambda})$
$=\inf_{u\in\Lambda}I(u)\,\inf_{\lambda>0}I(u_{\lambda})\left\{\lambda^{\alpha p-(d-sp)}\frac{1}{p}[u]^{p}_{W^{s,p}(\mathbb{R}^{d})}+\lambda^{-(d-\alpha q)}\frac{1}{q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}\right\}$

### v3

@@ 1. Introduction
(1.1) $\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{p^{2}}\log\left[C_{p}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx\right],$
@@
(1.2) $C_{p}=\frac{p}{d}\left(\frac{p-1}{e}\right)^{p-1}\pi^{-\frac{p}{2}}\left[\frac{\Gamma\left(\frac{d}{2}+1\right)}{\Gamma\left(d\frac{p-1}{p}+1\right)}\right]^{\frac{p}{d}}$
@@
(1.3) $\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{sp^{2}}\log\left(\mathcal{C}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy\right),$
@@
(1.4) $\mathcal{C}=C(d,p)\left(\frac{s(1-s)}{(d-sp)^{p-1}}\right),$
@@
(1.5) $\lim_{s\to 1}(1-s)\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy=K\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx.$
@@
A version of the above theorem on stratified Lie groups was proved by Ghosh, Kumar, and Ruzhansky in [15, Section $7$]. They established subelliptic fractional logarithmic Sobolev inequalities on stratified Lie groups with explicit constants depending on an additional parameter lying between $p$ and the fractional critical Sobolev exponent, under the condition $sp<d$. In contrast, Theorem 1.1 provides a constant $\mathcal{C}$ in Euclidean space that is independent of any extra parameter and also recovers the classical Sobolev logarithmic inequality.
@@
(1.6) $\int_{\mathbb{R}^{d}}|u(x)|^{2}\log\left(\frac{|u(x)|^{2}}{\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}}\right)\,dx+d(1+\log c)\|u\|^{2}_{L^{2}(\mathbb{R}^{d})}\leq\frac{c^{2}}{\pi}\int_{\mathbb{R}^{d}}|\nabla u(x)|^{2}\,dx.$
@@
(1.7) $\leq\frac{de^{p-1}c^{p}\,\mathcal{C}}{sp}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy,$
@@
(1.8) $\leq\frac{de^{p-1}c^{p}\,K}{p}\left(\frac{C(d,p)}{(d-p)^{p-1}}\right)\int_{\mathbb{R}^{d}}|\nabla u(x)|^{p}\,dx.$
@@
(1.9) $\int_{\mathbb{R}^{d}}|u(x)|^{p}\log|u(x)|\,dx\leq\frac{d}{(s-\alpha)p^{2}}\log\left(\mathcal{C}_{\alpha}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy\right),$
@@
(1.10) $\mathcal{D}^{p,q}_{s}=(W^{s,p}(\mathbb{R}^{d})\cap L^{q}(\mathbb{R}^{d}))\setminus\{0\}.$
@@
(1.11) $\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathscr{L}\,[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\,\|u\|^{1-a}_{L^{q}(\mathbb{R}^{d})},$
@@
(1.12) $a=\frac{d(q-p)}{(q-1)\bigl(dp-(d-sp)q\bigr)},$
@@ 2. Proof of fractional Sobolev logarithmic inequalities
(2.1) $\left(\int_{\mathbb{R}^{d}}|u(x)|^{p^{*}_{s}}\,dx\right)^{\frac{1}{p^{*}_{s}}}\leq\mathcal{C}^{\frac{1}{p}}\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy\right)^{\frac{1}{p}},$
@@
(2.2) $\mathcal{C}=C(d,p)\left(\frac{s(1-s)}{(d-sp)^{p-1}}\right),$
@@
(2.3) $\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathcal{C}^{\frac{a}{p}}[u]^{a}_{W^{s,p}(\mathbb{R}^{d})}\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a},$
@@
(2.5) $\int_{\mathbb{R}^{d}}|u(x)|^{q}\log\left(\frac{|u(x)|}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)\,dx\leq\frac{r}{r-q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$
@@
(2.6) $\log\|u\|_{L^{s}(\mathbb{R}^{d})}\leq\alpha(s)\log\|u\|_{L^{q}(\mathbb{R}^{d})}+(1-\alpha(s))\log\|u\|_{L^{r}(\mathbb{R}^{d})}.$
@@
(2.7) $s\Phi^{\prime}(s)=\frac{1}{\|u\|^{s}_{L^{s}(\mathbb{R}^{d})}}\int_{\mathbb{R}^{d}}|u(x)|^{s}\log\left(\frac{|u(x)|}{\|u\|_{L^{s}(\mathbb{R}^{d})}}\right)\,dx.$
@@
(2.8) $=\left(\frac{qr}{s^{2}(r-q)}\right)\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$
@@
(2.9) $\frac{1}{a}\log\left(\frac{\|u\|_{L^{r}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right)-\log\left(\mathcal{C}^{\frac{1}{p}}\right)\leq\log\left(\frac{[u]_{W^{s,p}(\mathbb{R}^{d})}}{\|u\|_{L^{q}(\mathbb{R}^{d})}}\right).$
@@
(2.10) $\frac{sp}{d}\int_{\mathbb{R}^{d}}\frac{|u(x)|^{p}}{\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}}\log\left(\frac{|u(x)|}{\|u\|_{L^{p}(\mathbb{R}^{d})}}\right)\,dx-\log\left(\mathcal{C}^{\frac{1}{p}}\right)\leq\log\left(\frac{[u]_{W^{s,p}(\mathbb{R}^{d})}}{\|u\|_{L^{p}(\mathbb{R}^{d})}}\right).$
@@ 3. Proof of Weighted fractional Sobolev logarithmic inequality
(3.1) $\|u\|_{W^{s,p,\alpha}(\mathbb{R}^{d})}=\left(\|u\|^{p}_{L^{p}(\mathbb{R}^{d})}+[u]^{p}_{W^{s,p,\alpha}(\mathbb{R}^{d})}\right)^{\frac{1}{p}}.$
@@
(3.2) $\||x|^{\gamma}u\|_{L^{\tau}(\mathbb{R}^{d})}\leq C[u]^{a}_{W^{s,p,\alpha}(\mathbb{R}^{d})}\||x|^{\beta}u\|^{(1-a)}_{L^{q}(\mathbb{R}^{d})},\hskip 8.5359pt\forall\ u\in C^{1}_{c}(\mathbb{R}^{d}).$
@@
(3.3) $\left(\int_{\mathbb{R}^{d}}|u(x)|^{\tau}\,dx\right)^{\!\frac{1}{\tau}}\leq C\left(\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy\right)^{\!\frac{1}{p}},$
@@
(3.4) $\forall\ u\in W^{s,p,\alpha}(\mathbb{R}^{d}).$
@@
(3.5) $\mathcal{D}^{p,q}_{s,\alpha}=(W^{s,p,\alpha}(\mathbb{R}^{d})\cap L^{q}(\mathbb{R}^{d}))\setminus\{0\}.$
@@
(3.6) $\|u\|_{L^{r}(\mathbb{R}^{d})}\leq\mathcal{C}_{\alpha}^{\frac{a}{p}}[u]_{W^{s,p,\alpha}(\mathbb{R}^{d})}^{a}\|u\|_{L^{q}(\mathbb{R}^{d})}^{1-a},$
@@
(3.8) $\leq\frac{d\,e^{p-1}c^{p}\,\mathcal{C}_{\alpha}}{(s-\alpha)p}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}|x|^{\alpha_{1}p}|y|^{\alpha_{2}p}\,dx\,dy.$
@@
(4.1) $\inf_{s>0}f(s)=f\left(\left(\frac{bN}{aM}\right)^{\frac{1}{a+b}}\right)=\frac{a+b}{a}\left(\frac{b}{a}\right)^{\frac{-b}{a+b}}M^{\frac{b}{a+b}}N^{\frac{a}{a+b}}.$
@@
(4.2) $I(u):=\frac{1}{p}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{|u(x)-u(y)|^{p}}{|x-y|^{d+sp}}\,dx\,dy+\frac{1}{q}\int_{\mathbb{R}^{d}}|u(x)|^{q}\,dx.$
@@
It is easy to deduce that for any $u\in\Lambda$, we have $u_{\lambda}\in\Lambda$ for all $\lambda>0$. Consequently, $\inf_{\lambda>0}I(u_{\lambda})\leq I(u_{1})=I(u)$, which implies
(4.4) $\inf_{u\in\Lambda}\inf_{\lambda>0}I(u_{\lambda})\leq\inf_{u\in\Lambda}I(u).$
Now set $a=\alpha p-(d-sp)>0$, $b=d-\alpha q>0$, $M=\frac{1}{p}[u]^{p}_{W^{s,p}(\mathbb{R}^{d})}>0$, and $N=\frac{1}{q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}>0$. By Lemma 4.1, there exists a unique $\lambda_{u}>0$ such that $\inf_{\lambda>0}I(u_{\lambda})=I(u_{\lambda_{u}})$ for every $u\in\Lambda$. Therefore, using the fact that for any $u\in\Lambda$ we have $u_{\lambda_{u}}\in\Lambda$, we obtain
$\inf_{u\in\Lambda}I(u)\leq I(u_{\lambda_{u}})=\inf_{\lambda>0}I(u_{\lambda}),\quad\forall\,u\in\Lambda\quad\Rightarrow\quad\inf_{u\in\Lambda}I(u)\leq\inf_{u\in\Lambda}\inf_{\lambda>0}I(u_{\lambda}).$
Applying Lemma 4.1 once again with the above parameters, and using the two estimates obtained above, we conclude that
$\inf_{u\in\Lambda}I(u)$ $=\inf_{u\in\Lambda}\,\inf_{\lambda>0}I(u_{\lambda})$
$=\inf_{u\in\Lambda}\,\inf_{\lambda>0}\left\{\lambda^{\alpha p-(d-sp)}\frac{1}{p}[u]^{p}_{W^{s,p}(\mathbb{R}^{d})}+\lambda^{-(d-\alpha q)}\frac{1}{q}\|u\|^{q}_{L^{q}(\mathbb{R}^{d})}\right\}$